CONSIDER A CIRCUIT HAVING 30 VOLT SOURCE , SWITCH , RESISTANCE R , CAPACITOR 0.01 MICRO FARAD AND RESISTANCE 10 KILO OHM CONNECTED IN SERIES WHEN THE SWITCH IN THIS CIRCUIT IS CLOSED , THE VOLTAGE ACROSS THE CAPACITOR MUST RISE TO 6.59 VOLT IN 0.1 MS IDENTIFY THE VALUE OF R THAT SHOULD BE USED


1
Expert's answer
2021-01-05T12:17:19-0500

Solution:

"v=v_o[1-e^{(\tfrac{-t}{\tau})}]"


6.59 = 30[1–"e^{(\tfrac{-0.0001}{\tau})}"]


1–"e^{(\tfrac{-0.0001}{\tau})}=0.2197"


"e^{(\tfrac{-0.0001}{\tau})}=0.7803"


"\dfrac{-0.0001}{\tau}=ln(0.7803)=-0.2481"


"\tau=\dfrac{0.0001}{0.2481}=0.0004031\;s"


RC = τ = 0.0004031

R = "\dfrac{0.0004031}{0.01e^{-6}}=40310\;\Omega"


added R = 40310 – 10000 = 30310 Ω

the voltage on a cap, charging

"v=v_o[1-e^{(\tfrac{-t}{\tau})}]"

v₀ is the battery voltage

v is the voltage after time t

R is resistance in ohms,

C is capacitance in farads

t is time in seconds

RC = τ = time constant